OpenStudy (anonymous):

express in simplest form : cubed root of a^6b^9/-64 show work :)

OpenStudy (anonymous):

cube root of a^6 is a^2 becuase a^2 x a^2 x a^2 = a^6 cube root of b^9 is b^3 becuase b^3 x b^3 x b^3 = b^9 cube root of -64 is -4 becuase (-4)(-4)(-4) = - 64 so the expression simplifies to (a^2b^3)/-4

OpenStudy (anonymous):

so would (a^2b^3)/-4 that be my final answer?